OpenStudy (anonymous):

Possible negative, positive, and complex numbers? f(x) = –2x3 – 5x2 – 6x + 4

OpenStudy (maheshmeghwal9):

?

OpenStudy (anonymous):

I meant zeros... my bad

OpenStudy (maheshmeghwal9):

if A, B, C are zeroes then for a cubic polynomial; \[x^3+(A+B+C)x^2+(AB+BC+CA)x+ABC=0.\] now u can put the values:)

OpenStudy (maheshmeghwal9):

A+B+C=5/2 AB+BC+CA=6/2=3 ABC=-4/2=-2. NOW JUST FIND:) A, B, C.

OpenStudy (maheshmeghwal9):

Got it or not? Plz tell when u understand this? & Ask doubt if u have^_^

OpenStudy (anonymous):

not... :s

OpenStudy (maheshmeghwal9):

where is ur doubt?

OpenStudy (anonymous):

x3+(A+B+C)x2+(AB+BC+CA)x+ABC=0 .... I dont get this part..

OpenStudy (maheshmeghwal9):

A+B+C IS x^2's coefficent ^ so on

OpenStudy (maheshmeghwal9):

it is a form of representing cubic polynomial just as in quadratic \[x^2-(A+B)+AB=0.\] if A & B are its zeroes

OpenStudy (anonymous):

question... is it 2 or 0 positive... 1 negative... and 0 or 2 complex?

OpenStudy (maheshmeghwal9):

sorry i can't do that; I don't know how to that:) but i gave u what I knew

OpenStudy (anonymous):

ok thanks

OpenStudy (maheshmeghwal9):

:)